OpenStudy (anonymous):

Please help! ill give medal Select the inequality that corresponds to the given graph. number line with a closed circle on negative 2 and 5 with shading in between

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

|2x - 4| > 2 |4x + 2| > 10 |4x - 6| <=14 |6x + 4| <=16

OpenStudy (anonymous):

do you know how?

OpenStudy (anonymous):

|2x - 4| > 2 2x - 4 > -2 or 2x - 4 > 2 2x > -2 + 4 2x > 2 + 4 2x > 2 2x > 6 x > 1 x > 3 not the first one |4x + 2| > 10 4x + 2 > -10 or 4x + 2 > 10 4x > -10-2 4x > 10 - 2 4x > -12 4x > 8 x > -3 x > 2 not the second one |4x - 6| <=14 -14 <=4x - 6<=14 -14 + 6 <= 4x - 6 + 6 <= 14 + 6 -8 <= 4x <= 20 -8/4 <= x <= 20/4 -2 <= x <= 5 Here is your answer.....

OpenStudy (anonymous):

omg thank you!!! can you do oone mor The number line below represents which combined inequality? number line with a closed circle on -6 and 5 with shading in between

OpenStudy (anonymous):

You can avoid having to try A and B first if you recognize that this must have an equal sign in the answer (<= or >=) since the circle is solid

OpenStudy (anonymous):

so true......I feel stupid....I didn't even have to check the first two

OpenStudy (anonymous):

The number line below represents which combined inequality? number line with a closed circle on -6 and 5 with shading in between x =>-6 or x <=5 x => -6 and x <= 5 x <= -6 and x => 5 x <= -6 or x => 5

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

its A...wait...A and B are the same answers

OpenStudy (anonymous):

no ones and, and ones or

OpenStudy (anonymous):

oh....your right....it is B

OpenStudy (anonymous):

ok thanks and one more Consider the absolute value inequality |x + 4| <=6. Part 1: Using complete sentences, explain whether this inequality will be an "and" compound inequality (conjunction) or an "or" compound inequality (disjunction) and why. Part 2: Provide the solution to the inequality and describe the graph in complete sentences

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

|x + 4| <= 6 hahahaha perfect.......check it out http://www.algebra.com/algebra/homework/absolute-value/absolute-value.faq.question.629030.html

OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3). @kelliegirl33

OpenStudy (anonymous):

Is that really an x2 or did you mean \[x ^{2}\]

OpenStudy (anonymous):

x squared

OpenStudy (anonymous):

OK. That's easy. Just plug in 3 for the x in the equation. \[f(3) = 3^{2}-1=9-1=8\]

OpenStudy (anonymous):

thank you! @BangkokGarrett Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x less than or greater to 12 A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right. A number line with an open circle on -1, a closed circle on 4, and shading in between. A number line with a closed circle on -1, shading to the left, and an open circle on 4, shading to the right. A number line with a closed circle on -1, an open circle on 4, and shading in between.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

x - 2 > -3 3x <= 12 x > -3 + 2 x <= 12/3 x > -1 x <= 4 2nd choice

OpenStudy (anonymous):

The depth d(x), in feet, of a circular swimming pool at a point x feet from the edge is represented by the function below. The function d of x equals 1 plus one-eighth x] If the depth varies from 1 foot to 4 feet, which inequality describes the domain of the function @kelliegirl33

OpenStudy (anonymous):

I do not know this one...sorry

OpenStudy (anonymous):

During a science experiment, Faye recorded the temperature of a solution as it cooled in a controlled environment for 6 minutes. She graphed the change in temperature as shown below. Straight line graph in the first quadrant titled Cooling Solution, shows Time (in minutes) on the x axis and Temperature (in degree centigrade) on the y axis. The x axis scale is from 0 to 9 in increments of 1. The y axis scale is from 0 to 45 in increments of 5. The graph intercepts the y axis at 40, passes through (3, 35), and ends at (6, 30). If the solution continues to cool at the same rate, how many minutes will it take for the temperature to cool to 0°C @kelliegirl33

OpenStudy (anonymous):

BangkokGarrett...you take this one....I am not good at these

OpenStudy (anonymous):

24 19 14 9 answer chocies

OpenStudy (anonymous):

On the depth problem 1 + 1/8 x has to be between 1 and for. Solve 1 + 1/8x >= 1 and you get x >=0 Solve 1 + 1/8x <=4 and you get x <=24. So x is x is between 0 and 24 feet. 0 <= x <= 24

OpenStudy (anonymous):

ok so its 24 ?

OpenStudy (anonymous):

@BangkokGarrett

OpenStudy (anonymous):

Ooops 4, not for. what a weird typo. hehe

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

@BangkokGarrett

OpenStudy (anonymous):

What is it? 24?